# Cell-Free Massive MIMO with Hardware-Impaired Wireless Fronthaul thanks:

### v1

@@ IV Spectral Efficiency with Optimal Linear Combiner
$\mathbf{R}_{k}$ $=\kappa_{\mathrm{frt}}\kappa_{\mathrm{ac}}\sum_{i\neq k}p_{i}\sum_{l=1}^{L}\mathbf{b}_{il}\mathbf{b}_{il}^{\mbox{\tiny$\mathrm{H}$}}$
$\quad+\kappa_{\mathrm{frt}}\sum_{l=1}^{L}\mathbf{G}_{l}\mathbf{P}_{l}(\mathbf{D}_{\mathrm{ac},l}+\sigma^{2}\mathbf{I}_{N})\mathbf{P}_{l}^{\mbox{\tiny$\mathrm{H}$}}\mathbf{G}_{l}^{\mbox{\tiny$\mathrm{H}$}}$ (15)
@@ VI Numerical Results
In Fig. 4, we compare two scenarios: (i) only access-side impairments, with $\kappa_{\rm frt}=1$ and $\kappa_{\rm ac}=0.9$, and (ii) both access- and fronthaul-side impairments, with $\kappa_{\rm frt}=\kappa_{\rm ac}=0.9$. As illustrated in the figure, when the fronthaul is also impaired, even distortion-aware combining cannot surpass the performance of the distortion-unaware scheme in the case with no fronthaul impairments. For both scenarios, distortion-aware processing yields higher SE, particularly for high-rate UEs. However, for mid-rate UEs, distortion-awareness becomes significantly more critical when both access and fronthaul chains are impaired, compared to the case where only the access side is impaired.

### v2

@@ IV Spectral Efficiency with Optimal Linear Combiner
$\mathbf{R}_{k}$ $=\kappa_{\mathrm{frt}}\kappa_{\mathrm{ac}}\sum_{i\neq k}p_{i}\left(\sum_{l=1}^{L}\mathbf{b}_{il}\right)\left(\sum_{l^{\prime}=1}^{L}\mathbf{b}_{il^{\prime}}^{\mbox{\tiny$\mathrm{H}$}}\right)$
$\quad+\kappa_{\mathrm{frt}}\sum_{l=1}^{L}\mathbf{G}_{l}\mathbf{P}_{l}(\mathbf{D}_{\mathrm{ac},l}+\sigma^{2}\mathbf{I}_{N})\mathbf{P}_{l}^{\mbox{\tiny$\mathrm{H}$}}\mathbf{G}_{l}^{\mbox{\tiny$\mathrm{H}$}}$
@@ VI Numerical Results
In Fig. 4, we compare two scenarios: (i) only access-side impairments, with $\kappa_{\rm frt}=1$ and $\kappa_{\rm ac}=0.9$, and (ii) both access- and fronthaul-side impairments, with $\kappa_{\rm frt}=\kappa_{\rm ac}=0.9$. As illustrated in the figure, when the fronthaul is also impaired, even distortion-aware combining cannot surpass the performance of the distortion-unaware scheme in the case with no fronthaul impairments. For both scenarios, distortion-aware processing yields higher SE, particularly for high-rate UEs. However, for mid-rate UEs, distortion-awareness becomes more critical when both access and fronthaul chains are impaired, compared to the case where only the access side is impaired.